# Quantum Adaptive Self-Attention for Quantum Transformer Models

### v2

@@ 4 Results and Discussions
Component $QASA_{classical}$ $QASA$
@@
Encoder Layers 4$\times$ Transformer (4 heads) 3$\times$ Transformer + 1$\times$ QuantumEncoderLayer
@@ 4.1 Experiment Details
  • $QASA$: A quantum-enhanced hybrid transformer in which the final transformer encoder block is replaced with a quantum encoder layer that incorporates a variational quantum circuit (VQC). The VQC uses 8 qubits and 4 layers, with RX/RZ rotations and entangling CNOT gates, and outputs expectation values of Pauli-Z measurements. These values are projected back into the hidden dimension space and processed in a residual feed-forward fashion.
@@ 4.2 Results
To compare model performance, we report the validation metrics at epoch 45 for all three architectures: Vanilla Transformer, $QASA_{classical}$, and $QASA$. The results are summarized in the table below:
Table 3: Validation performance comparison at final epoch.
Model val_mse val_mae
Transformer 0.5188 0.3946
$QASA_{classical}$ 0.0122 0.0916
$QASA$ 0.0085 0.0679
From the results in Table LABEL:tab:mae_mse, it is evident that the proposed $QASA$ model outperforms both baselines across all two metrics. Specifically, $QASA$ achieves the lowest mean squared error (MSE: 0.0085), and mean absolute error (MAE: 0.0679).
The $QASA_{classical}$ model ranks second, with a MSE of 0.0122 and a MAE of 0.0916. Meanwhile, the standard Transformer exhibits the weakest performance, yielding a significantly higher MSE of 0.5188 and MAE of 0.3946.
Quantitatively, $QASA$ achieves approximately a 30% reduction in MSE compared to its classical counterpart, and an impressive 98% reduction relative to the vanilla Transformer. These results highlight the substantial advantage of incorporating quantum layers in temporal regression tasks.
Table 4: Performance Comparison Between Classical and Quantum Models
Task Model MAE MSE MAE Std MSE Std
ARMA Classical 1.2812 2.6682 0.0865 0.3964
Quantum 0.9611 1.4735 0.0639 0.2631
Chaotic Logistic Classical 0.4294 0.2247 0.0422 0.0471
Quantum 0.2942 0.1169 0.0322 0.0236
Damped Oscillator Classical 0.1100 0.0132 0.0248 0.0051
Quantum 0.3024 0.1105 0.1476 0.0869
Noisy Damped Osc Classical 0.0273 0.0011 0.0032 0.0003
Quantum 0.1035 0.0162 0.0554 0.0157
Piecewise Regime Classical 11.7206 137.4312 0.2486 5.7949
Quantum 11.7273 137.8328 0.6064 14.1222
Sawtooth Classical 0.2279 0.0526 0.0278 0.0119
Quantum 0.4844 0.2519 0.044 0.032
Square Wave Classical 0.8731 1.2323 0.0675 0.0792
Quantum 0.9145 1.2643 0.0786 0.2427
Seasonal Trend Classical 0.5477 0.3262 0.0478 0.0539
Quantum 0.6448 0.4443 0.0938 0.1210
To evaluate the effectiveness of the proposed Quantum Adaptive Self-Attention (QASA) architecture, we conducted experiments across eight synthetic time-series forecasting tasks, each characterized by different signal dynamics—ranging from periodicity and chaos to noise and discontinuities. The result in Table LABEL:tab:ts_mae_mse Performance was assessed using Mean Absolute Error (MAE) and Mean Squared Error (MSE), along with their respective standard deviations, comparing QASA with a classical Transformer baseline under identical training settings.
In the ARMA task, which exhibits autoregressive temporal dependencies, QASA achieved substantially lower error metrics (MAE: 0.9611, MSE: 1.4735) compared to the classical model (MAE: 1.2812, MSE: 2.6682), indicating the quantum model’s superior capability in capturing structured, yet noisy, linear dynamics. Similarly, in the chaotic logistic map—a highly nonlinear and sensitive system—QASA outperformed its classical counterpart by a notable margin (MAE: 0.2942 vs. 0.4294; MSE: 0.1169 vs. 0.2247), suggesting that the parameterized quantum circuits were effective in modeling long-range chaotic correlations.
However, the performance gains did not generalize uniformly across all tasks. In the damped oscillator and noisy damped oscillator tasks, the classical model significantly outperformed QASA, achieving both lower error and more stable results. For instance, in the noisy damped oscillator scenario, the classical model reached a near-zero MSE of 0.0011, while QASA yielded a substantially higher MSE of 0.0162. These results suggest that QASA may be more sensitive to noise in periodic systems or that its entangled representation does not always translate well in low-frequency, low-variance regimes.
In the piecewise regime task, which includes abrupt temporal discontinuities, both models performed similarly in terms of absolute error (MAE: 11.72 for both), but the quantum model demonstrated higher variability (MSE Std: 14.1222 vs. 5.7949), pointing to instability in handling sharp transitions. Similarly, in the sawtooth and square wave tasks, which feature high-frequency discontinuities, the classical model outperformed QASA with a considerable margin, suggesting that the quantum-enhanced attention mechanism struggles with sharply segmented or discontinuous signals.
The seasonal trend task, characterized by smooth periodicity and trend components, also saw the classical model outperform QASA (MAE: 0.5477 vs. 0.6448), further reinforcing the conclusion that conventional attention mechanisms may be better suited for capturing predictable harmonic patterns without the overhead of quantum encoding.
Across all tasks, it is notable that the QASA model exhibited higher variance in both MAE and MSE, pointing to less stable training dynamics. This behavior may stem from the sensitivity of variational quantum circuits to parameter initialization and gradient noise. Nevertheless, in tasks where nonlinear temporal dependencies dominate, such as ARMA and chaotic logistic systems, QASA consistently outperforms the classical baseline, demonstrating the potential of hybrid quantum-classical models to learn expressive, high-dimensional temporal representations.
Overall, these findings highlight both the promise and the limitations of quantum-enhanced attention. While QASA introduces meaningful performance improvements in certain classes of time-series tasks—particularly those characterized by chaos or autoregressive behavior—it currently lacks the robustness and generality required to replace classical architectures across the board. Future work may address these gaps by refining the quantum encoding strategies, incorporating hybrid training objectives, or developing adaptive layer selection mechanisms that better match quantum resources to signal complexity.
5 Quantum Adaptive Self-Attention: Complexity Analysis
We propose a quantum adaptive self-attention mechanism that replaces parts of the classical attention module—specifically the softmax and matrix multiplication components—with parameterized quantum circuits (PQC), forming a hybrid quantum-classical architecture. This modification allows us to bypass some known classical complexity limitations in gradient computation.
Recent work in fine-grained complexity theory has shown that, under the Strong Exponential Time Hypothesis (SETH), the gradient computation of the classical transformer attention cannot be faster than $O(T^{2})$ [AS24] (Here $T$ can be viewed as a size of the matrix, see equation 8). However, it is well known that SETH does not hold in the quantum regime due to Grover’s algorithm [Gro96], which provides a quadratic speedup for unstructured search problems. Grover’s algorithm solves $k$-SAT in $O(\sqrt{2}^{n})=O(1.414^{n})$ time (here $n$ is the number of variables), violating the classical SETH assumption. This speedup is also known to be optimal, as established by the so-called BBBV bound [BBBV97]..
This observation aligns with our numerical findings, where we observe that the hybrid quantum attention mechanism appears more efficient in practice. While a rigorous proof is currently out of reach due to the hybrid nature of the model, this empirical evidence supports the hypothesis that quantum mechanisms can outperform classical counterparts in this context. We leave a formal theoretical investigation of this potential quantum advantage to future work.
5.1 Background: Classical SETH and Gradient Lower Bounds

### v3

@@ 4 Results and Discussions
Component $_{classical} QASAQASA$
@@
[…]
@@ 4.1 Experiment Details
  • QASAQASA: A quantum-enhanced hybrid transformer in which the final transformer encoder block is replaced with a quantum encoder layer that incorporates a variational quantum circuit (VQC). The VQC uses 8 qubits and 4 layers, with $R_{X}$ and $R_{Z}$ rotations and entangling CNOT gates, and outputs expectation values of Pauli-$Z$ measurements. These values are projected back into the hidden dimension space and processed in a residual feed-forward fashion.
@@
For the broader benchmark evaluation (Table 5), we use a reduced hidden dimension of 64 (with 4 attention heads and 4 encoder layers) to ensure tractable quantum circuit simulation across all nine tasks. Training is conducted for 200 epochs using the AdamW optimizer (learning rate $5\times 10^{-4}$, weight decay $10^{-4}$) with cosine annealing scheduling. Each configuration is evaluated over five random seeds to assess consistency. Under this configuration, the QASA model has 201,405 total parameters and the classical Transformer baseline has 200,257 parameters—a difference of less than 0.6%, ensuring a fair parameter-matched comparison.
Computational Cost.
All experiments are conducted on a single CPU (Apple M-series) using PennyLane’s lightning.qubit simulator for the quantum circuit components. Table 4 reports the wall-clock training time per task (200 epochs) averaged across the nine benchmark tasks. The classical Transformer completes training in approximately 4.5 minutes per task (1.4 s/epoch), while QASA requires approximately 2 hours and 7 minutes per task (40 s/epoch)—roughly a $\mathbf{29\times}$ slowdown. This overhead is entirely attributable to quantum circuit simulation: each forward pass through the QuantumLayer requires executing the parameterized quantum circuit for every token in the sequence, with gradient computation via the parameter-shift rule doubling the number of circuit evaluations. On actual quantum hardware, circuit execution time scales with circuit depth ($L_{q}\times$ gate time) rather than with the exponential cost of classical state-vector simulation, and the overhead would be substantially reduced.
Table 4: Average wall-clock training time per task (200 epochs, single CPU, hidden_dim=64). The $29\times$ overhead arises from simulating the 9-qubit PQC classically; on quantum hardware, circuit execution would be substantially faster.
Model Time/task sec/epoch
Classical Transformer 4.5 min 1.4 s
QASA 2h 07m 40.4 s
Slowdown factor $\approx 29\times$
4.2 Dataset Construction
To probe the task-conditional behaviour of quantum-enhanced attention, we construct nine univariate synthetic time-series benchmarks, each designed to expose a distinct signal regime—smooth periodic, chaotic, stochastic, trend-dominated, discontinuous, and mixed. All sequences contain $T=500$ points and are split chronologically into 80% training and 20% testing. For all tasks, the model is trained to predict $x_{t+1}$ from a sliding context window of length $L=20$. Real-world validation is performed on the ETTh1 dataset; see the corresponding paragraph below.
Synthetic tasks.
Let $t\in\{0,1,\dots,T{-}1\}$ denote the discrete time index and $\tau=50\,t/(T{-}1)$ the normalised continuous time on $[0,50]$. The nine signals are defined as follows:
  1. 1.
    ARMA: A second-order autoregressive moving-average process,
    $x_{t}=0.75\,x_{t-1}-0.25\,x_{t-2}+\varepsilon_{t}+0.65\,\varepsilon_{t-1},\qquad\varepsilon_{t}\sim\mathcal{N}(0,1).$ (36)
  2. 2.
    Chaotic logistic map: A deterministic chaotic system,
    $x_{t}=r\,x_{t-1}(1-x_{t-1}),\qquad r=3.9,\ x_{0}=0.2.$ (37)
    The parameter $r=3.9$ places the system deep in the chaotic regime.
  3. 3.
    Damped oscillator: A clean decaying sinusoid,
    $x(\tau)=e^{-0.05\,\tau}\sin(\tau).$ (38)
  4. 4.
    Noisy damped oscillator: As above with additive Gaussian noise,
    $x_{t}=e^{-0.05\,t}\sin(2\pi\cdot 0.2\,t)+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,0.05^{2}).$ (39)
  5. 5.
    Piecewise regime: A concatenation of five linear regimes with two impulsive shocks at $t\in[200,205)$ and $t\in[350,355)$, mimicking structural breaks,
    $x_{t}=\alpha_{k}\,t+\beta_{k}+\eta_{t},\quad\eta_{t}\sim\mathcal{N}(0,0.01^{2}),\ (\alpha_{k},\beta_{k})\text{ regime-dependent}.$ (40)
  6. 6.
    Sawtooth wave: A periodic ramp with period $P=10$,
    $x(\tau)=2\left(\frac{\tau}{P}-\left\lfloor\tfrac{1}{2}+\frac{\tau}{P}\right\rfloor\right).$ (41)
  7. 7.
    Square wave: A binary waveform with time-varying period and duty cycle,
    $x(\tau)=\mathrm{sgn}\!\left(d(\tau)-\phi(\tau)\right),$ (42)
    where $\phi(\tau)=(\tau\bmod P(\tau))/P(\tau)$, $P(\tau)=5+5\sin(0.1\tau)$, and $d(\tau)=0.2+0.3\left(1+\sin(0.05\tau)\right)$.
  8. 8.
    Seasonal trend with noise: An additive mixture of trend, seasonality, and noise,
    $x_{t}=0.03\,t+\sin(0.4\,t)+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,0.2^{2}).$ (43)
  9. 9.
    Waveform: A smooth reference sinusoid, $x(\tau)=\sin(\tau)$.
These nine tasks span four qualitatively different dynamical regimes: (i) stochastic but stationary (ARMA), (ii) chaotic (logistic map), (iii) smooth periodic/quasi-periodic (damped oscillator, waveform, sawtooth, square wave), and (iv) non-stationary or discontinuous (piecewise regime, seasonal trend, noisy damped oscillator). This coverage lets us identify when quantum enhancement helps rather than making a single aggregate claim.
Real-world dataset (ETTh1).
The Electricity Transformer Temperature dataset [40] provides hourly measurements of oil temperature from an electricity transformer over approximately two years. We use a subsampled training set of 500 windows (sequence length $L=20$) with the target variable “OT” (oil temperature), and evaluate on a held-out test set of 500 windows. No dataset-specific hyperparameter tuning is performed; all models use the same configuration as the synthetic benchmarks.
@@ 4.3 Results
Initial experiments with a larger model capacity (hidden_dim=256, 45 epochs) suggested that QASA could outperform classical models on specific tasks, motivating the systematic parameter-matched comparison in Table 5.